# FreeAskWorld: An Interactive and Closed-Loop Simulator for Human-Centric Embodied AI

### v1

@@ Introduction
  • An interactive LLM-driven simulation framework for dynamic, realistic human agents with high-level intention modeling and semantic interaction control
@@
Core behavioral challenges in social navigation span from individual-level proxemics to macro-level social signaling in multi-agent contexts (Rios-Martinez, Spalanzani, and Laugier 2015; Mavrogiannis et al. 2023). Traditional simulation platforms have primarily relied on low-level physical behavior models to simulate crowd dynamics and agent-level reactions to robots (Grzeskowiak et al. 2021; Tsoi et al. 2022; Vuong et al. 2024; Pérez-Higueras et al. 2023). While effective for modeling local interactions and physical constraints, these approaches often miss the dynamic and contextual nuances of societal-level human behavior. To address this, recent work has turned to generative AI to simulate high-level, context-aware behaviors(Wang et al. 2024b; Xi et al. 2025; Piao et al. 2025). For instance, MARPLE introduces a hierarchical structure that decomposes behavior into missions, subgoals, and atomic actions (Jin et al. 2024). Recent simulators, such as Virtual Community(Zhou et al. 2025), MetaUrban(Wu et al. 2024), and Grutopia(Wang et al. 2024a), push this further to urban-scale world by modeling open-ended social interactions among autonomous generative agents.
@@
Wayfinding through verbal directions is a complex and dynamic process, shaped by the direction giver, the environment, the task at hand, and the recipient(Hund, Haney, and Seanor 2008). Two common strategies in verbal instructions are the route perspective, using egocentric cues like landmarks and left or right turns; and the survey perspective, based on map-like descriptions involving distances, street names, and cardinal directions. Navigation style from the direction giver’s perspective is primarily influenced by socio-spatial background and gender (Galea and Kimura 1993; Lawton 1996, 2001; Kato and Takeuchi 2003). For example, irregular European layouts tend to favor landmark-based navigation, while grid-based American cities rely more on street names(Hund, Schmettow, and Noordzij 2012). Women usually use more spatial references, provide longer instructions, and incorporate hedging expressions(Sing and Kalingga 2011).
@@
The classical VLN task was introduced by the Room-to-Room benchmark (Anderson et al. 2018), which provides natural language navigation instructions grounded in real-world indoor panoramic scenes. However, R2R relies on discrete navigation graphs, limiting realism and fine-grained spatial interaction. REVERIE (Qi et al. 2020) extends R2R by introducing object grounding, but still lacks continuous motion and physical embodiment. R2R-CE (Krantz et al. 2020) addresses this by enabling continuous action spaces with real-time physics, introducing challenges like low-level control and precise localization. Talk2Nav (Vasudevan, Dai, and Van Gool 2021) further advances VLN with long-range instruction following using dual attention mechanisms and spatial memory modules.
Despite progress, VLN systems face three key limitations: (1) reliance on static, one-shot instructions, reducing adaptability to dynamic environments and multi-turn interactions; (2) lack of integration between high-level planning and social intention modeling, limiting agents’ ability to interpret social cues and perform context-aware behaviors; and (3) simulators that fail to capture the complexity of real-world environments with dynamic human agents, moving vehicles, and socially grounded interactions (Wang et al. 2024a).
@@
The visual and kinematic realism of human agents significantly impacts the credibility of Sim2Real navigation systems. Many navigation simulators still rely on static 3D character models with minimal animation diversity(Li et al. 2021; Vuong et al. 2024). HA-VLN takes a step forward by combining SMPL-based body models with AI-based action planning, enhancing the naturalness of character motion (Dong et al. 2025).
@@
To support consistent embodiment and context-sensitive behaviors, we adopt MotionX(Lin et al. 2023) as the SMPL-X animation library and structure its motions into high-level categories and subcategories, allowing the selection of semantically relevant motions for each activity. We also developed a custom SMPL-X animation driver plugin that allows seamless access to the full range of animations within the library. To enhance animation fidelity and character realism, we integrated Blending functionality into the motion controller, enabling smooth transitions between actions.
@@
Method 1: We propose a new framework that uses Multimodal Large Language Models (MLLMs) to create diverse virtual human appearances controlled by natural language. By combining UV mapping with semantic profiles (such as gender, occupation, and ethnicity), we build a comprehensive pipeline for texture generation guided by language, enabling scalable and consistent material creation across a variety of scenarios. At the mesh level, we introduce variation by randomly modifying the shape parameters of the SMPL-X model, producing diverse body types characterized by differences in height, weight, and proportions. This approach, coupled with semantic control over beta shapes properties, enhances both the realism and expressiveness of human simulations. Method 2: We leverage the Synbody (Yang et al. 2023) dataset to generate SMPL-X models, which incorporate clothing, footwear, hair, and other detailed features, offering a more immersive and realistic virtual experience. This approach enhances the fidelity of modeling human appearance.
@@
  • Success weighted by Path Length (SPL): Measures both success and efficiency (Anderson et al. 2018), defined as $\mathrm{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{l_{i}^{*}}{\max(l_{i},l_{i}^{*})}$, where $S_{i}\in\{0,1\}$ indicates success in episode $i$, $l_{i}^{*}$ is the shortest path length to the goal, and $l_{i}$ is the actual path length taken by the agent.
@@
We compare our dataset with other existing VLN datasets, including Talk2Nav (Vasudevan, Dai, and Van Gool 2021), R2R (Anderson et al. 2018), REVERIE (Qi et al. 2020), ScaleVLN (Wang et al. 2023), NavRAG (Wang et al. 2025), GSA-R2R (Li et al. 2024a), HA-R2R (Li et al. 2024a), NaVid (Zhang et al. 2024), DynamicVLN (Sun, Qiu, and Aoki 2025), and VLN-Video (Li et al. 2024b). Table 1 summarizes their characteristics.
@@
We used Unity Perception(Borkman et al. 2021) to construct a rich and diverse synthetic dataset that encompasses multiple types of annotations and data modalities. The dataset is designed to support a wide range of vision and navigation tasks and includes both dense per-frame annotations and global scene-level metadata. Specifically, the dataset comprises:
  • 2D and 3D bounding boxes: for object localization in image and world space.
  • Instance segmentation masks: for object-level understanding.
  • Semantic segmentation maps: for category-level scene parsing.
  • Surface normal maps: to provide geometric orientation information.
  • Depth maps: capture of scene geometry and spatial structure.
  • Panoramic RGB images and six 90 ° perspective images: for full environmental perception.
  • Human-Computer Interaction data: including natural language instructions, dialog histories, and agent trajectories.
  • 2D occupancy heatmaps: representing spatial accessibility and the distribution of static objects within the environment, thereby facilitating downstream tasks such as mapping and localization.
  • Environment metadata: encompassing map-level specifications such as spatial boundaries, semantic region definitions, and other contextual information about the virtual environment.
As shown in Figure 4, the six main types of synthetic data are briefly summarized.
@@
Based on the 2D occupancy heatmaps that encode the layout of static elements, along with the 3D bounding boxes that capture the positions of dynamic entities, the simulation environment can be accurately reconstructed. By further integrating the provided world information, it becomes possible to generate a comprehensive digital twin of the scene. This reconstructed environment enables open-loop evaluations similar to those in the nuScenes dataset (Caesar et al. 2020), and is particularly well suited for unstructured settings as in FreeAD (Peng et al. 2025), and supports a broad spectrum of downstream tasks including navigation planning, behavior prediction, and human-computer interaction studies.
@@
  • ETPNav (An et al. 2024): A hierarchical VLN‑CE framework that performs online topological mapping, cross-modal planning with transformers, and low-level control using a rotate‑then‑forward schema augmented by obstacle‑avoidance heuristics.
  • BEVBert (An et al. 2022): A map‑based multimodal pre-training model that leverages hybrid topo‑metric representations to improve spatial reasoning and language-guided navigation robustness.
@@ Results
In closed-loop experiments, the human baseline demonstrates that agents capable of querying for additional navigation instructions significantly improve pathfinding accuracy (from 40.2% to 82.6% ). This improvement is attributed to scene complexity, such as identical stores appearing in different directions within 50 meters (shown in Fig. 4), where humans may occasionally experience shallow memory of directions and become disoriented. Fine-tuned models, ETPNav-FT and BEVBert-FT, show substantial improvements over their base counterparts in both Navigation Error and Oracle Navigation Error. The increase in Trajectory Length suggests enhanced familiarity with the scene and broader exploration. Although ETPNav-FT achieves partial destination success, overcoming the zero Oracle Success Rate observed in baseline models, its overall Success Rate remains zero. A similar trend is observed in the Social Mobile Manipulation task of InfiniteWorld (Ren et al. 2024), where tasks involving social interactions substantially reduce robot performance. This challenge arises from weak dynamic social navigation, pedestrian/vehicle collisions, and limitations in long-range planning, abstract reasoning, memory retention, and higher-level decision-making—areas that warrant further investigation. Additionally, BEVBert consistently outperforms ETPNav across both evaluation metrics, reinforcing its state-of-the-art performance in Vision-and-Language Navigation (VLN) and highlighting commonalities with our task.
@@
Human (no asking) 47.5 40.2 38.2 18.3 41.3 11.3 0.0
Human (asking) 59.9 82.6 71.2 3.49 82.6 1.63 0.78

### v2

@@ Introduction
  • An interactive, LLM-driven simulation framework featuring dynamic and realistic human agents with high-level intention modeling, semantic interaction control, and a synthetic data generation pipeline.
@@
Core behavioral challenges in social navigation span from individual-level proxemics to macro-level social signaling in multi-agent contexts (rios2015proxemics; 10.1145/3583741). Traditional simulation platforms have primarily relied on low-level physical behavior models to simulate crowd dynamics and agent-level reactions to robots (grzeskowiak2021crowd; tsoi2022sean; vuong2024habicrowd; perez2023hunavsim). While effective for modeling local interactions and physical constraints, these approaches often miss the dynamic and contextual nuances of societal-level human behavior. To address this, recent work has turned to generative AI to simulate high-level, context-aware behaviors(wang2024survey; xi2025rise; piao2025agentsociety). For instance, MARPLE introduces a hierarchical structure that decomposes behavior into missions, subgoals, and atomic actions (jin2024marple). Recent simulators, such as Virtual Community(zhou2025virtual), MetaUrban(wu2024metaurban), and Grutopia(wang2024grutopia), push this further to urban-scale world by modeling open-ended social interactions among autonomous generative agents.
@@
Wayfinding through verbal directions is a complex and dynamic process, shaped by the direction giver, the environment, the task at hand, and the recipient(hund2008role). Two common strategies in verbal instructions are the route perspective, using egocentric cues like landmarks and left or right turns; and the survey perspective, based on map-like descriptions involving distances, street names, and cardinal directions. Navigation style from the direction giver’s perspective is primarily influenced by socio-spatial background and gender (galea1993sex; lawton1996strategies; lawton2001gender; kato2003individual). For example, irregular European layouts tend to favor landmark-based navigation, while grid-based American cities rely more on street names(hund2012impact). Women usually use more spatial references, provide longer instructions, and incorporate hedging expressions(sing2011gender).
@@
The classical VLN task was introduced by the Room-to-Room benchmark (anderson2018vision), which provides natural language navigation instructions grounded in real-world indoor panoramic scenes. However, R2R relies on discrete navigation graphs, limiting realism and fine-grained spatial interaction. REVERIE (qi2020reverie) extends R2R by introducing object grounding, but still lacks continuous motion and physical embodiment. R2R-CE (krantz2020beyond) addresses this by enabling continuous action spaces with real-time physics, introducing challenges like low-level control and precise localization. Talk2Nav (vasudevan2021talk2nav) further advances VLN with long-range instruction following using dual attention mechanisms and spatial memory modules.
Despite progress, VLN systems face three key limitations: (1) reliance on static, one-shot instructions, reducing adaptability to dynamic environments and multi-turn interactions; (2) lack of integration between high-level planning and social intention modeling, limiting agents’ ability to interpret social cues and perform context-aware behaviors; and (3) simulators that fail to capture the complexity of real-world environments with dynamic human agents, moving vehicles, and socially grounded interactions (wang2024grutopia).
@@
The visual and kinematic realism of human agents significantly impacts the credibility of Sim2Real navigation systems. Many navigation simulators still rely on static 3D character models with minimal animation diversity(li2021igibson; vuong2024habicrowd). HA-VLN takes a step forward by combining SMPL-based body models with AI-based action planning, enhancing the naturalness of character motion (dong2025ha).
@@
To support consistent embodiment and context-sensitive behaviors, we adopt MotionX(lin2023motionx) as the SMPL-X animation library and structure its motions into high-level categories and subcategories, allowing the selection of semantically relevant motions for each activity. We also developed a custom SMPL-X animation driver plugin that allows seamless access to the full range of animations within the library. To enhance animation fidelity and character realism, we integrated Blending functionality into the motion controller, enabling smooth transitions between actions.
@@
Method 1: We propose a new framework that uses Multimodal Large Language Models (MLLMs) to create diverse virtual human appearances controlled by natural language. By combining UV mapping with semantic profiles (such as gender, occupation, and ethnicity), we build a comprehensive pipeline for texture generation guided by language, enabling scalable and consistent material creation across a variety of scenarios. At the mesh level, we introduce variation by randomly modifying the shape parameters of the SMPL-X model, producing diverse body types characterized by differences in height, weight, and proportions. This approach, coupled with semantic control over beta shapes properties, enhances both the realism and expressiveness of human simulations. Method 2: We leverage the Synbody (yang2023synbody) dataset to generate SMPL-X models, which incorporate clothing, footwear, hair, and other detailed features, offering a more immersive and realistic virtual experience. This approach enhances the fidelity of modeling human appearance.
@@
  • Success weighted by Path Length (SPL): Measures both success and efficiency (anderson2018vision), defined as $\mathrm{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{l_{i}^{*}}{\max(l_{i},l_{i}^{*})}$, where $S_{i}\in\{0,1\}$ indicates success in episode $i$, $l_{i}^{*}$ is the shortest path length to the goal, and $l_{i}$ is the actual path length taken by the agent.
@@
We compare our dataset with other existing VLN datasets, including Talk2Nav (vasudevan2021talk2nav), R2R (anderson2018vision), REVERIE (qi2020reverie), ScaleVLN (wang2023scaling), NavRAG (wang2025navrag), GSA-R2R (li2024human), HA-R2R (li2024human), NaVid (zhang2024navid), DynamicVLN (sun2025dynamicvln), and VLN-Video (li2024vln). Table 1 summarizes their characteristics.
@@
We used Unity Perception(borkman2021unity) to construct a rich and diverse synthetic dataset that encompasses multiple types of annotations and data modalities. The dataset is designed to support a wide range of vision and navigation tasks and includes both dense per-frame annotations and global scene-level metadata. Specifically, the dataset comprises:
  • Visual annotations: 2D/3D bounding boxes, instance and semantic segmentation.
  • Geometric annotations: depth maps and surface normal maps for scene geometry.
  • Visual observations: panoramic RGB images and six 90 perspective views.
  • Interaction data: natural language instructions, dialog histories, and agent trajectories.
  • Spatial representations: 2D occupancy heatmaps for mapping and localization.
  • Environment metadata: map boundaries, semantic regions, and other contextual information.
As shown in Figure 4, the six main types of synthetic data are summarized.
@@
Based on the 2D occupancy heatmaps that encode the layout of static elements, along with the 3D bounding boxes that capture the positions of dynamic entities, the simulation environment can be accurately reconstructed. By further integrating the provided world information, it becomes possible to generate a comprehensive digital twin of the scene. This reconstructed environment enables open-loop evaluations similar to those in the nuScenes dataset (caesar2020nuscenes), and is particularly well suited for unstructured settings as in FreeAD (peng2025bench2freead), and supports a broad spectrum of downstream tasks including navigation planning, behavior prediction, and human-computer interaction studies.
@@
  • ETPNav (an2024etpnav): A hierarchical VLN‑CE framework that performs online topological mapping, cross-modal planning with transformers, and low-level control using a rotate‑then‑forward schema augmented by obstacle‑avoidance heuristics.
  • BEVBert (an2022bevbert): A map‑based multimodal pre-training model that leverages hybrid topo‑metric representations to improve spatial reasoning and language-guided navigation robustness.
@@ Results
In closed-loop experiments, the human baseline demonstrates that agents capable of querying for additional navigation instructions significantly improve pathfinding accuracy (from 40.2% to 82.6% ). This improvement is attributed to scene complexity, such as identical stores appearing in different directions within 50 meters (shown in Fig. 4), where humans may occasionally experience shallow memory of directions and become disoriented. Fine-tuned models, ETPNav-FT and BEVBert-FT, show substantial improvements over their base counterparts in both Navigation Error and Oracle Navigation Error. The increase in Trajectory Length suggests enhanced familiarity with the scene and broader exploration. Although ETPNav-FT achieves partial destination success, overcoming the zero Oracle Success Rate observed in baseline models, its overall Success Rate remains zero. A similar trend is observed in the Social Mobile Manipulation task of InfiniteWorld (ren2024infiniteworld), where tasks involving social interactions substantially reduce robot performance. This challenge arises from weak dynamic social navigation, pedestrian/vehicle collisions, and limitations in long-range planning, abstract reasoning, memory retention, and higher-level decision-making—areas that warrant further investigation. Additionally, BEVBert consistently outperforms ETPNav across both evaluation metrics, reinforcing its state-of-the-art performance in Vision-and-Language Navigation (VLN) and highlighting commonalities with our task.
@@
Human(no ask) 47.5 40.2 38.2 18.3 41.3 11.3 0.0
Human(ask) 59.9 82.6 71.2 3.49 82.6 1.63 0.78